# SFT Doesn’t Always Hurt General Capabilities: Revisiting Domain-Specific Fine-Tuning in LLMs

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of general-purpose tasks, including question answering, mathematical reasoning, and code generation (DBLP:journals/corr/abs-2412-15115; yang2025qwen3; touvron2023llama; dubey2024llama). To further enhance their effectiveness in specialized applications, practitioners often perform additional supervised fine-tuning (SFT) using domain-specific data. This process enriches the model with domain knowledge and yields substantial performance gains on domain-specific tasks (labrak2024biomistral; lin2024panacea; peng2024ecellm). SFT has thus become a standard paradigm for adapting LLMs to real-world deployment scenarios.
However, recent studies have shown that fine-tuning LLMs on domain-specific datasets can substantially impair their generalization capabilities (huan2025does; lin2025rec; chen2025sft; bansal2025context; sanyal2025upweighting; chu2025sft; shenfeld2025rl). For example, performing SFT on LLMs like Qwen-3 (yang2025qwen3) or Gemma-3 (team2025gemma) using domain-specific datasets, such as those from e-commercial or biomedical domains, often leads to significant performance drops on general-purpose benchmarks such as GSM8K (cobbe2021training), HumanEval (chen2021evaluating), or IFEval (zhou2023instruction), which assess core capabilities like mathematical reasoning, code generation, and instruction following. This phenomenon raises the need for a closer examination of domain-specific SFT.
@@
The first observation is relatively expected, since smaller learning rates naturally suppress parameter drift compared to more aggressive updates (pareja2025unveiling). The second, however, is more surprising. Prior to the LLM era, practical experience in machine and deep learning suggested that larger learning rates are often essential for better downstream performance (mohtashami2023special; li2019towards; sadrtdinov2024large). In contrast, we show that LLMs behave differently: comparable domain-specific performance can still be achieved under smaller learning rates. We further provide a theoretical analysis supporting this phenomenon. In addition, a closer inspection of prior studies shows that their strong degradation claims predominantly arise under relatively large learning rates. Taken together, our empirical and theoretical evidence demonstrates that a careful choice of learning rate offers a practical path to balance domain adaptation with general capability preservation (§3).
While adopting smaller learning rates typically yields a better trade-off, we also observe that this does not fully mitigate the general-performance degradation in all cases. To address this, we further investigate mitigation approaches that could mitigate such degradation (§4). Specifically, we assess a range of representative strategies evaluated in sanyal2025upweighting, including L2 regularization, LoRA (hu2022lora), model averaging (wortsman2022robust), and FLOW (sanyal2025upweighting), along with our proposed method, Token-Adaptive Loss Reweighting (TALR). TALR adaptively down-weights hard tokens by solving a constrained optimization problem that admits a closed-form solution, thereby tempering their potential disproportionate influence on general capability degradation during training. Our experiments demonstrate that TALR provides advantages in further suppressing general-performance degradation compared to these baselines. Nevertheless, no existing method including TALR can completely eliminate the degradation, highlighting the need for more advanced strategies to be explored in future work.
@@ 2 Related Work
Our problem setting can be broadly framed within the scope of continual learning, where models must acquire new knowledge while retaining previously learned capabilities to avoid catastrophic forgetting. Existing approaches are typically divided into two categories: data-dependent and data-oblivious. Data-dependent methods assume access to a subset of the training data from earlier stages, whereas data-oblivious methods rely solely on the pre-trained model without revisiting any prior data. The latter is particularly realistic in practice, as access to proprietary or large-scale pre-training corpora is often infeasible, yet it remains relatively underexplored. For a broader overview of this landscape, we refer readers to recent surveys (wang2024comprehensive). Our focus in this paper is on the data-oblivious setting.
Data-oblivious approaches. One line of work introduces loss regularization to constrain the fine-tuned model from drifting too far from its initialization, such as L2 regularization in parameter space (kumar2025maintaining; kirkpatrick2017overcoming). Another line of work explores the idea of model averaging, which combines the parameters of the pre-trained model and the fully fine-tuned model through a convex combination, aiming to balance adaptation with retention (wortsman2022robust; lubana2022quadratic; ilharco2023editing; kleiman2025soup). LoRA (hu2022lora; biderman2024lora) represents another widely used strategy, enforcing low-rank updates to the weight matrices so that parameter changes are confined to a restricted subspace, thereby limiting catastrophic drift while improving efficiency. Besides, data reweighting has been explored as a promising strategy; for example, FLOW (sanyal2025upweighting) mitigates forgetting in vision tasks by adjusting the loss weights of easy and hard samples.
Extensions to LLMs. Existing research has primarily focused on data-dependent methods in the LLM context (scialom2022fine; yin2023dynosaur; wang2024inscl; xiong2023rationale; mok2023large). In contrast, data-oblivious approaches remain relatively underexplored, though several studies have begun adapting ideas from continual learning on traditional models to LLMs (sanyal2025upweighting; razdaibiedina2023progressive; wang2023orthogonal; zhao2024sapt). Refer to the survey by wu2024continuallearninglargelanguage for more details. However, LLMs differ substantially from earlier architectures in scale, pre-training regimes, and emergent capabilities, which makes their adaptation dynamics distinct. As a result, we revisit continual SFT of LLMs on domain-specific datasets, aiming to better understand its mechanism.
@@
In this section, we revisit the role of learning rate in domain-specific SFT and its impact on general capability degradation. Surprisingly, we find that using a smaller learning rate (e.g., $1\mathrm{e}{-6}$) can substantially reduce the loss of general capabilities, while achieving domain-specific task performance on par with much larger learning rates. This suggests that the severe degradation reported in prior work may stem, at least in part, from overly aggressive optimization (huan2025does; lin2025rec; chen2025sft; bansal2025context; sanyal2025upweighting; shenfeld2025rl). Indeed, many of these studies used relatively large learning rates such as $5\mathrm{e}{-6}$ or $2\mathrm{e}{-5}$.
To systematically investigate this effect, we experiment on two domain-specific datasets: MedCalc (khandekar2024medcalc) and ESCI (reddy2022shopping). We choose these datasets because existing open-source LLMs perform poorly on them, making them representative scenarios where domain-specific SFT is most motivated: to enhance specialized capabilities in domains where the initialized model is weak. Below are details of the experimental setups and results for each dataset.
@@ 3.1.1 Datasets
MedCalc (khandekar2024medcalc) consists of 10.1k training and 1.05k test examples. Each instance includes a brief patient note and a clinical instruction (e.g., “What is the patient’s CHA2DS2-VASc score?”), with the goal of predicting a numeric, categorical, or datetime answer. The training set provides gold chain-of-thought (CoT) rationales, which we use as supervision targets during SFT. The ESCI dataset (reddy2022shopping) is an e-commerce product classification benchmark containing query–product pairs labeled as Exact, Substitute, Complement, or Irrelevant. The training set consists of 49k examples, and the test set contains 10k examples. We consider two training settings: w/ CoT, where the target sequence includes both reasoning and the final label, and w/o CoT, where it contains only the label.
@@ 3.1.2 Evaluation Protocol
For the MedCalc task, we follow the evaluation protocol of khandekar2024medcalc and report accuracy based on the model’s final answer. For general capability evaluation, we measure performance on a suite of general-purpose benchmarks using the lm-evaluation-harness framework (eval-harness), following the same evaluation setup as prior works (lin2025rec; sanyal2025upweighting; bansal2025context). Model checkpoints are selected based on their best performance on the target domain task, after which the corresponding models are evaluated on general-purpose benchmarks, reflecting practical scenarios where downstream task performance is prioritized (sanyal2025upweighting; bansal2025context). The evaluation metric for each benchmark is detailed in Appendix C.1. Since ESCI is highly imbalanced across classes, we follow prior work on imbalanced classification (xu2024whole; xu2025pisces) and report balanced accuracy (BACC) as our primary metric.
@@ 3.3 Theoretical Analysis
To better understand the empirical phenomena observed previously, we provide a theoretical analysis from the perspective of information theory. Motivated by the equivalence between language modeling and data compression (deletang2024language; DBLP:conf/acl/JiWQ0Z0LDL025), we view an LLM as a compressor, where the effectiveness of training can be measured through changes in code length. In this view, improvements or degradations in performance across datasets correspond to variations in compression rate. Below, we formalize this perspective by introducing the notion of token trees and describing the LLM compression protocol in our context.
@@
Having established the compression protocol, we now follow prior work (deletang2024language; DBLP:conf/acl/JiWQ0Z0LDL025) and use changes in expected code length as a surrogate metric for an LLM’s modeling quality on a given dataset distribution. In this view, reductions in code length discrepancy correspond to better alignment with the data distribution. Formally, this is captured by the following proposition.
@@ 4.2 Results
We evaluate all mitigation strategies considered in sanyal2025upweighting, including L2 regularization, LoRA, Wise-FT (model averaging), and FLOW, together with our proposed TALR. Table 1 and 2 reports the trade-off between domain performance and general performance under two learning rates. The baseline configurations follow sanyal2025upweighting.
@@ 5.2 Broader Impacts
Better domain adaptation. Our findings provide practitioners with insights when developing domain-specific LLMs. Taking the medical domain as an example, jeong2024limited show that existing medical-specialized LLMs often fail to outperform their corresponding initialized LLMs. This suggests that the quality of domain-specific data alone may not be as high as in the sophisticated post-training pipelines applied to base models. Hence, methods that preserve as much of the initialized LLM’s general capabilities as possible while injecting domain knowledge may lead to stronger overall performance.
Mitigating exploration loss in SFT warm-up for RLVR. Before reinforcement learning with verifiable reward (RLVR), SFT is often used as a warm-up step to inject knowledge (lin2025training) or align formats (guo2025deepseek). However, excessive SFT can over-stabilize the model, causing its output trajectories to become rigid and thereby undermining exploration during RL training. In contrast, models prior to excessive SFT typically exhibit more diverse behaviors. Thus, strategies that mitigate general-performance degradation and preserve the base model’s diversity may help alleviate this issue and enable more effective RL.
Contents of Appendix
@@ B.1 Preliminaries
Lossless Compression. In lossless compression, the goal is to encode a sequence of symbols $x=(x_{1},x_{2},\dots,x_{T})$ drawn from a source distribution $P$ into a binary representation without any loss of information, such that the original sequence can be perfectly reconstructed. According to Shannon’s source coding theorem (shannon1948mathematical), the limit of compression is given by the Shannon entropy of the source: $H(P):=\mathbb{E}_{x\sim P}\big[-\log P(x)\big]$, which specifies the minimum expected number of bits per symbol needed for encoding.
LLM Modeling is Compression. Given a dataset $\mathcal{D}$ drawn from the true distribution $P$ and a model distribution $Q$, the expected code length under arithmetic coding (witten1987arithmetic) is given by $H(P,Q):=\mathbb{E}_{x\sim P}\big[-\log Q(x)\big]$. Thus, minimizing the log-likelihood loss directly corresponds to reducing the expected compression rate when the model is employed as a lossless compressor (deletang2024language; DBLP:series/txtcs/Hutter05; DBLP:conf/acl/JiWQ0Z0LDL025).
@@ B.2 LLM Compression Protocol
Our goal is to analyze the dynamics of domain-specific SFT. Motivated by the equivalence between language modeling and data compression (deletang2024language; DBLP:conf/acl/JiWQ0Z0LDL025), we view an LLM as a compressor, where the effectiveness of training can be measured through changes in code length. In this view, improvements or degradations in performance across datasets correspond to variations in compression rate. Below, we formalize this perspective by introducing the notion of token trees and describing the LLM compression protocol in our context.
@@
Based on the above, we adopt the expected code length as a surrogate metric for an LLM’s modeling quality on a given dataset (deletang2024language). Specifically, reductions in code length discrepancy indicate better alignment between the model distribution and the data distribution, whereas increases suggest deterioration. This perspective will serve as the foundation for our subsequent analysis.
@@ B.3 Approximating Fine-tuning Dynamics with Exponential Tilting
Our goal in this part is to uncover the behavior of LLMs during domain-specific fine-tuning. Fine-tuning alters the conditional distributions assigned to each node in the token tree, thereby shifting the model’s alignment with the data distribution. For analytical simplicity, we view fine-tuning at a high level as introducing perturbations to the probability assigned to each token node. To approximate these dynamics, we adopt the lens of exponential tilting (maity2023understanding), which captures how the distribution is reweighted under incremental updates.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of general-purpose tasks, including question answering, mathematical reasoning, and code generation (Yang et al., 2024; 2025; Touvron et al., 2023; Dubey et al., 2024). To further enhance their effectiveness in specialized applications, practitioners often perform additional supervised fine-tuning (SFT) using domain-specific data. This process enriches the model with domain knowledge and yields substantial performance gains on domain-specific tasks (Labrak et al., 2024; Lin et al., 2024; Peng et al., 2024). SFT has thus become a standard paradigm for adapting LLMs to real-world deployment scenarios.
However, recent studies have shown that fine-tuning LLMs on domain-specific datasets can substantially impair their generalization capabilities (Huan et al., 2025; Lin et al., 2025a; Chen et al., 2025; Bansal and Sanghavi, 2025; Sanyal et al., 2025; Chu et al., 2025; Shenfeld et al., 2025). For example, performing SFT on LLMs like Qwen-3 (Yang et al., 2025) or Gemma-3 (Team et al., 2025) using domain-specific datasets, such as those from e-commercial or biomedical domains, often leads to significant performance drops on general-purpose benchmarks such as GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), or IFEval (Zhou et al., 2023), which assess core capabilities like mathematical reasoning, code generation, and instruction following. This phenomenon raises the need for a closer examination of domain-specific SFT.
@@
The first observation is relatively expected, since smaller learning rates naturally suppress parameter drift compared to more aggressive updates (Pareja et al., 2025). The second, however, is more surprising. Prior to the LLM era, practical experience in machine and deep learning suggested that larger learning rates are often essential for better downstream performance (Mohtashami et al., 2023; Li et al., 2019; Sadrtdinov et al., 2024). In contrast, we show that LLMs behave differently: comparable domain-specific performance can still be achieved under smaller learning rates. We further provide a theoretical analysis supporting this phenomenon. In addition, a closer inspection of prior studies shows that their strong degradation claims predominantly arise under relatively large learning rates. Taken together, our empirical and theoretical evidence demonstrates that a careful choice of learning rate offers a practical path to balance domain adaptation with general capability preservation (§3).
While adopting smaller learning rates typically yields a better trade-off, we also observe that this does not fully mitigate the general-performance degradation in all cases. To address this, we further investigate mitigation approaches that could mitigate such degradation (§4). Specifically, we assess a range of representative strategies evaluated in Sanyal et al. (2025), including L2 regularization, LoRA (Hu et al., 2022), model averaging (Wortsman et al., 2022), and FLOW (Sanyal et al., 2025), along with our proposed method, Token-Adaptive Loss Reweighting (TALR). TALR adaptively down-weights hard tokens by solving a constrained optimization problem that admits a closed-form solution, thereby tempering their potential disproportionate influence on general capability degradation during training. Our experiments demonstrate that TALR provides advantages in further suppressing general-performance degradation compared to these baselines. Nevertheless, no existing method including TALR can completely eliminate the degradation, highlighting the need for more advanced strategies to be explored in future work.
@@ 2 Related Work
Our problem setting can be broadly framed within the scope of continual learning, where models must acquire new knowledge while retaining previously learned capabilities to avoid catastrophic forgetting. Existing approaches are typically divided into two categories: data-dependent and data-oblivious. Data-dependent methods assume access to a subset of the training data from earlier stages, whereas data-oblivious methods rely solely on the pre-trained model without revisiting any prior data. The latter is particularly realistic in practice, as access to proprietary or large-scale pre-training corpora is often infeasible, yet it remains relatively underexplored. For a broader overview of this landscape, we refer readers to recent surveys (Wang et al., 2024a). Our focus in this paper is on the data-oblivious setting.
Data-oblivious approaches. One line of work introduces loss regularization to constrain the fine-tuned model from drifting too far from its initialization, such as L2 regularization in parameter space (Kumar et al., 2025; Kirkpatrick et al., 2017). Another line of work explores the idea of model averaging, which combines the parameters of the pre-trained model and the fully fine-tuned model through a convex combination, aiming to balance adaptation with retention (Wortsman et al., 2022; Lubana et al., 2022; Ilharco et al., 2023; Kleiman et al., 2025). LoRA (Hu et al., 2022; Biderman et al., 2024) represents another widely used strategy, enforcing low-rank updates to the weight matrices so that parameter changes are confined to a restricted subspace, thereby limiting catastrophic drift while improving efficiency. Besides, data reweighting has been explored as a promising strategy; for example, FLOW (Sanyal et al., 2025) mitigates forgetting in vision tasks by adjusting the loss weights of easy and hard samples.
Extensions to LLMs. Existing research has primarily focused on data-dependent methods in the LLM context (Scialom et al., 2022; Yin et al., 2023; Wang et al., 2024b; Xiong et al., 2023; Mok et al., 2023). In contrast, data-oblivious approaches remain relatively underexplored, though several studies have begun adapting ideas from continual learning on traditional models to LLMs (Sanyal et al., 2025; Razdaibiedina et al., 2023; Wang et al., 2023; Zhao et al., 2024). Refer to the survey by Wu et al. (2024) for more details. However, LLMs differ substantially from earlier architectures in scale, pre-training regimes, and emergent capabilities, which makes their adaptation dynamics distinct. As a result, we revisit continual SFT of LLMs on domain-specific datasets, aiming to better understand its mechanism.
@@
In this section, we revisit the role of learning rate in domain-specific SFT and its impact on general capability degradation. Surprisingly, we find that using a smaller learning rate (e.g., $1\mathrm{e}{-6}$) can substantially reduce the loss of general capabilities, while achieving domain-specific task performance on par with much larger learning rates. This suggests that the severe degradation reported in prior work may stem, at least in part, from overly aggressive optimization (Huan et al., 2025; Lin et al., 2025a; Chen et al., 2025; Bansal and Sanghavi, 2025; Sanyal et al., 2025; Shenfeld et al., 2025). Indeed, many of these studies used relatively large learning rates such as $5\mathrm{e}{-6}$ or $2\mathrm{e}{-5}$.
To systematically investigate this effect, we experiment on two domain-specific datasets: MedCalc (Khandekar et al., 2024) and ESCI (Reddy et al., 2022). We choose these datasets because existing open-source LLMs perform poorly on them, making them representative scenarios where domain-specific SFT is most motivated: to enhance specialized capabilities in domains where the initialized model is weak. Below are details of the experimental setups and results for each dataset.
@@ 3.1.1 Datasets
MedCalc (Khandekar et al., 2024) consists of 10.1k training and 1.05k test examples. Each instance includes a brief patient note and a clinical instruction (e.g., “What is the patient’s CHA2DS2-VASc score?”), with the goal of predicting a numeric, categorical, or datetime answer. The training set provides gold chain-of-thought (CoT) rationales, which we use as supervision targets during SFT. The ESCI dataset (Reddy et al., 2022) is an e-commerce product classification benchmark containing query–product pairs labeled as Exact, Substitute, Complement, or Irrelevant. The training set consists of 49k examples, and the test set contains 10k examples. We consider two training settings: w/ CoT, where the target sequence includes both reasoning and the final label, and w/o CoT, where it contains only the label.
@@ 3.1.2 Evaluation Protocol
For the MedCalc task, we follow the evaluation protocol of Khandekar et al. (2024) and report accuracy based on the model’s final answer. For general capability evaluation, we measure performance on a suite of general-purpose benchmarks using the lm-evaluation-harness framework (Gao et al., 2024), following the same evaluation setup as prior works (Lin et al., 2025a; Sanyal et al., 2025; Bansal and Sanghavi, 2025). Model checkpoints are selected based on their best performance on the target domain task, after which the corresponding models are evaluated on general-purpose benchmarks, reflecting practical scenarios where downstream task performance is prioritized (Sanyal et al., 2025; Bansal and Sanghavi, 2025). The evaluation metric for each benchmark is detailed in Appendix C.1. Since ESCI is highly imbalanced across classes, we follow prior work on imbalanced classification (Xu et al., 2024; 2025) and report balanced accuracy (BACC) as our primary metric.
@@ 3.3 Theoretical Analysis
To better understand the empirical phenomena observed previously, we provide a theoretical analysis from the perspective of information theory. Motivated by the equivalence between language modeling and data compression (Deletang et al., 2024; Ji et al., 2025), we view an LLM as a compressor, where the effectiveness of training can be measured through changes in code length. In this view, improvements or degradations in performance across datasets correspond to variations in compression rate. Below, we formalize this perspective by introducing the notion of token trees and describing the LLM compression protocol in our context.
@@
Having established the compression protocol, we now follow prior work (Deletang et al., 2024; Ji et al., 2025) and use changes in expected code length as a surrogate metric for an LLM’s modeling quality on a given dataset distribution. In this view, reductions in code length discrepancy correspond to better alignment with the data distribution. Formally, this is captured by the following proposition.
@@ 4.2 Results
We evaluate all mitigation strategies considered in Sanyal et al. (2025), including L2 regularization, LoRA, Wise-FT (model averaging), and FLOW, together with our proposed TALR. Table 1 and 2 reports the trade-off between domain performance and general performance under two learning rates. The baseline configurations follow Sanyal et al. (2025).
@@ 4.3 Token-Level Analysis
Output: Updated parameters $\theta$
1ex
foreach training step do
   Token NLLs: $\ell_{t}\leftarrow-\log p_{t}$;
 Adaptive weights with lower-bound clipping:
Let $N$ be the number of supervised tokens in the batch;
 Mean (averaged) reweighted loss:
$\mathcal{L}_{\text{TALR}}\;=\;\frac{1}{N}\sum_{t=1}^{N}w_{t}\,(-\log p_{t})$
Parameter update:
$\theta\leftarrow\theta\;-\;\eta\,\nabla_{\theta}\mathcal{L}_{\text{TALR}}$
end foreach
@@ 5.2 Broader Impacts
Better domain adaptation. Our findings provide practitioners with insights when developing domain-specific LLMs. Taking the medical domain as an example, Jeong et al. (2024) show that existing medical-specialized LLMs often fail to outperform their corresponding initialized LLMs. This suggests that the quality of domain-specific data alone may not be as high as in the sophisticated post-training pipelines applied to base models. Hence, methods that preserve as much of the initialized LLM’s general capabilities as possible while injecting domain knowledge may lead to stronger overall performance.
Mitigating exploration loss in SFT warm-up for RLVR. Before reinforcement learning with verifiable reward (RLVR), SFT is often used as a warm-up step to inject knowledge (Lin et al., 2025b) or align formats (Guo et al., 2025). However, excessive SFT can over-stabilize the model, causing its output trajectories to become rigid and thereby undermining exploration during RL training. In contrast, models prior to excessive SFT typically exhibit more diverse behaviors. Thus, strategies that mitigate general-performance degradation and preserve the base model’s diversity may help alleviate this issue and enable more effective RL.
@@ B.1 Preliminaries
Lossless Compression. In lossless compression, the goal is to encode a sequence of symbols $x=(x_{1},x_{2},\dots,x_{T})$ drawn from a source distribution $P$ into a binary representation without any loss of information, such that the original sequence can be perfectly reconstructed. According to Shannon’s source coding theorem (Shannon, 1948), the limit of compression is given by the Shannon entropy of the source: $H(P):=\mathbb{E}_{x\sim P}\big[-\log P(x)\big]$, which specifies the minimum expected number of bits per symbol needed for encoding.
LLM Modeling is Compression. Given a dataset $\mathcal{D}$ drawn from the true distribution $P$ and a model distribution $Q$, the expected code length under arithmetic coding (Witten et al., 1987) is given by $H(P,Q):=\mathbb{E}_{x\sim P}\big[-\log Q(x)\big]$. Thus, minimizing the log-likelihood loss directly corresponds to reducing the expected compression rate when the model is employed as a lossless compressor (Deletang et al., 2024; Hutter, 2005; Ji et al., 2025).
@@ B.2 LLM Compression Protocol
Our goal is to analyze the dynamics of domain-specific SFT. Motivated by the equivalence between language modeling and data compression (Deletang et al., 2024; Ji et al., 2025), we view an LLM as a compressor, where the effectiveness of training can be measured through changes in code length. In this view, improvements or degradations in performance across datasets correspond to variations in compression rate. Below, we formalize this perspective by introducing the notion of token trees and describing the LLM compression protocol in our context.
@@
Based on the above, we adopt the expected code length as a surrogate metric for an LLM’s modeling quality on a given dataset (Deletang et al., 2024). Specifically, reductions in code length discrepancy indicate better alignment between the model distribution and the data distribution, whereas increases suggest deterioration. This perspective will serve as the foundation for our subsequent analysis.
@@ B.3 Approximating Fine-tuning Dynamics with Exponential Tilting
Our goal in this part is to uncover the behavior of LLMs during domain-specific fine-tuning. Fine-tuning alters the conditional distributions assigned to each node in the token tree, thereby shifting the model’s alignment with the data distribution. For analytical simplicity, we view fine-tuning at a high level as introducing perturbations to the probability assigned to each token node. To approximate these dynamics, we adopt the lens of exponential tilting (Maity et al., 2023), which captures how the distribution is reweighted under incremental updates.